write an equation in point-slope form for the line through the given point with the given slope.
(-10,-1); m=-1

Answer by ankor@dixie-net.com(22740)

write an equation in point-slope form for the line through the given point with the given slope.
(-10,-1); m=-1
:
the point/slope form: y - y1 = m(x - x1), where
m = -1
x1 = -10
y1 = -1
:
y - (-1) = -1(x - (-10))
minus a minus is a plus
y + 1 = -1(x + 10)
y + 1 = -x - 10
y = -x - 10 - 1
y = -x - 11
x + y = -11